My Notes A tennis player receives a shot with the ball (0.060 0 kg) traveling horizontally at 55.0 m/s and returns the shot with the ball traveling horizontally at 39.5 m/s in the opposite direction. (Assume the initial direction of the ball is in the −x direction.)

(a) What is the impulse delivered to the ball by the tennis racquet?___________________ î N · s

(b) What work does the racquet do on the ball?______________________ J

My Notes An estimated force–time curve for a baseball struck by a bat . Let Fmax = 19 000 N, ta = 1.5 ms, and tb = 3 ms.

From this curve, determine each of the following.

(a) the magnitude of the impulse delivered to the ball N · s

(b) the average force exerted on the ball kN Need My Notes

A railroad car of mass 2.51 104 kg is moving with a speed of 4.02 m/s. It collides and couples with three other coupled railroad cars, each of the same mass as the single car and moving in the same direction with an initial speed of 2.01 m/s.

(a) What is the speed of the four cars after the collision? (Round your answer to at least two decimal places.) m/s

(b) How much mechanical energy is lost in the collision? J
Answer #1

a tennis player....

a)

I = m ( v - u)

I = 0.06* ( 39.5 + 55)

I = 5.67 Ns

====

b)

W = 0.5 m ( v^2 - u^2) = 0.5* 0.06* ( 39.5^2 - 55^2)

W = - 43.94 J

=====

plz post other problem separately

Comment before rate in case any doubt, will reply for sure.. goodluck